Where does the answer of (32) come from in the following question?

Q: By how much will RGDP change if businesses increase their level of investment spending by $8 billion and the MPC in the economy is ¾?

A: 32
Answer #1

Answer

Yes.

Multiplier =1/(1-MPC)

=1/(1-(3/4))

=4

change in RGDP =change in investments * multiplier

=8*4

=$32 billion